# Vehicle Rebalancing Under Adherence Uncertainty

### v1

@@ IV-F Baseline Model
$\sum_{c\in\mathcal{C}}x_{cj}\leq\rho\cdot\mathbb{E}[\mathbf{d_{j}}]\hskip 2.84544pt\forall j\in\mathcal{R}$
$x_{cj}\cdot\big{(}\mathcal{T}_{cj}-|\mathcal{H}|\big{)}\leq 0\hskip 2.84544pt\forall c\in\mathcal{C},j\in\mathcal{R}.$
Both the models are run within the same environment, the only difference being the rebalancing model. The drivers are initialized with 20% confidence level. The simulation in run for a period of 7 days (168 time steps) after which the driver confidence converges as shown in Fig 6. At each time step, we extract the evaluation metrics: Allocation (30), driver profit (31), confidence (32), and the met demand (33). Then the two models are compared based on these metrics.
IV-G Model Performance
This section highlights key observations from the results in Table III, focusing on allocation efficiency, driver profit, met demand, and system confidence under optimistic, neutral, and pessimistic settings. The percentage improvement for each metric is calculated as the average improvement across all fleet sizes and scenarios. The proposed rebalancing model demonstrates superior performance across multiple fleet sizes and evaluation metrics compared to the baseline model under most scenarios as shown in the Fig. 9.
The proposed model consistently achieves higher allocation rates than the baseline across all fleet sizes, particularly in neutral and pessimistic scenarios. For instance, at a fleet size of 4000 under pessimistic conditions, the proposed model achieves an allocation rate of 0.709, outperforming the baseline’s 0.661. Similarly, for a fleet size of 6000, the model achieves 0.634 compared to the baseline’s 0.518. On average, the proposed model improves allocation efficiency by 6.3% compared to the baseline.
Driver profit is significantly higher for the proposed model across all fleet sizes and scenarios. Notably, in the optimistic scenario for a fleet size of 2000, the proposed model yields the hourly driver profit of 11.242 compared to 10.065 for the baseline. This trend continues for larger fleets, with the proposed model maintaining higher profitability for all three classes of the drivers. The profit differences become more pronounced in case of the pessimistic drivers. For instance, at a fleet size of 8000 under pessimistic conditions, the proposed model achieves a profit of 6.737, significantly outperforming the baseline’s 5.347. The overall improvement in driver profit is 11.8% more as compared to the baseline.
The proposed model outperforms the baseline in meeting demand across all fleet sizes, especially under pessimistic scenarios where operational challenges of moving the drivers are most severe. At a fleet size of 6000 under pessimistic conditions, the proposed model meets 90.4% of the demand, compared to 74.0% for the baseline. Even in the optimistic scenario for smaller fleets (e.g., 2000), the proposed model performs slightly better than the baseline with a demand satisfaction rate of 0.374 compared to 0.373. On average, the proposed model improves demand satisfaction by 4.5%.
The confidence metric indicates system reliability and driver trust. While both models converge to a confidence level of 100% in optimistic and neutral scenarios, the proposed model demonstrates resilience under pessimistic conditions. For example, at a fleet size of 4000 under pessimistic conditions, the confidence score for the proposed model remains high at 0.940, significantly outperforming the baseline’s 0.470. For larger fleets like 8000, the baseline’s confidence drops dramatically to 0.030, whereas the proposed model maintains a reasonable level of 0.360. On average, the proposed model improves confidence by 14.7% under pessimistic conditions.

### v2

@@ IV-D Analysis of Adherence Aware Vehicle Rebalancing Problem
$\sum_{c\in\mathcal{C}}\mathbf{x_{cj}}\cdot\left(1-\frac{\mathcal{T}_{cj}}{\mathcal{H}}\right)\leq\rho\cdot\mathbf{\nu_{j}}\hskip 14.22636pt\forall j\in\mathcal{R}$
$\mathbf{x_{cj}}\cdot\left(\mathcal{T}_{cj}-\mathcal{H}\right)\leq 0\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The modeling assumptions of the above defined baseline [7] match exactly the proposed model, making it directly implementable. This model uses a parameter $\rho$ to control over-saturation of a region with taxis by bounding the total number of taxis recommended to a region. The authors in [7] do not discuss how to set the value of this parameter. We set $\rho=1$ to use the predicted demand as the target. The model’s objective function is designed to favor regions with higher demands.
Problem IV.3.
Baseline Model 2 (B2)
$\min_{\mathbf{x}}$ $\left\|\frac{\sum_{c}\mathbf{x_{cj}}}{N}-\frac{\mathbf{\nu_{j}}}{\sum_{j}\mathbf{\nu_{j}}}\right\|_{1}-\beta\sum_{c\in\mathcal{C}}\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}\cdot\mathcal{D}_{cj}$ (30)
s.t. $\sum_{j\in\mathcal{R}}\mathbf{x_{cj}}=1\hskip 2.84544pt\forall c\in\mathcal{C}$
$\mathbf{x_{cj}}\cdot\left(\mathcal{T}_{cj}-\mathcal{H}\right)\leq 0\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
The above model is a single-step finite horizon variant of the original one [6], which does not consider the travel time constraint. To avoid the scenario that a driver is recommended a destination which is not reachable within the planning horizon, a travel time constraint was included in IV.3.
Problem IV.4.
Baseline Model 3 (B3)
$\min_{\mathbf{X}}$ $\sum_{i}\sum_{j}\tau_{ij}\cdot\mathbf{X_{ij}}$ (31)
s.t. $V_{i}+\sum_{j}\mathbf{X_{ij}}-\sum_{j}\mathbf{X_{ij}}\geq\frac{\mathbf{\nu_{i}}}{\sum_{j}\mathbf{\nu_{j}}}\cdot N\hskip 14.22636pt\forall i\in\mathcal{R}$
$\sum_{j}\mathbf{X_{ij}}\leq V_{i}\hskip 14.22636pt\forall i\in\mathcal{R}$
The above model represents the feedback and proportional predictive rebalancing strategy [3]. The term on the right-hand side of the first constraint is the desired vehicle distribution. Note that the original model does not explicitly include the second constraint, but we include it in this work to ensure that the total number of taxis recommended to reposition in other regions of a region $i\in\mathcal{R}$ must be no more than the taxis that stand in that region. This model has a limitation: it does not consider the travel time feasibility in the case of finite planning horizon. As a result, not all the taxis dispatched to a region reach within the planning horizon. We could not find a direct way to add a constraint to limit such flows because adding a constraint leads to infeasibility due to the first constraint; because it is not always possible to achieve a proportional distribution under restricted flows.
Problem IV.5.
Baseline Model 4 (B4)
$\min_{\mathbf{X},\mathbf{Y},\mathbf{T},\mathbf{S}}$ $\sum_{i}\sum_{j}\mathcal{D}_{ij}\cdot\mathbf{X_{ij}}+\beta\sum_{i}\sum_{j}\mathcal{D}_{ji}\cdot\mathbf{Y_{ij}}+\gamma\sum_{i}\mathbf{T_{i}}$ (32)
s.t. $\sum_{j}\mathbf{Y_{ji}}\leq\mathbf{S_{i}},\sum_{j}\mathbf{Y_{ij}}\leq\mathbf{\nu_{i}},\sum_{j}\mathbf{X_{ij}}\leq V_{i}\hskip 8.5359pt\forall i\in\mathcal{R}$
$\mathbf{T_{i}}=\mathbf{\nu_{i}}-\sum_{j}Y_{ij}\hskip 14.22636pt\forall i\in\mathcal{R}$
$\mathbf{S_{i}}=V_{i}+\sum_{j}\mathbf{X_{ij}}-\sum_{j}\mathbf{X_{ij}}\hskip 14.22636pt\forall i\in\mathcal{R}$
$a_{ij}\cdot\mathbf{X_{ij}}=0,\hskip 14.22636ptb_{ij}\cdot\mathbf{Y_{ij}}=0\hskip 14.22636pt\forall i\in\mathcal{R},j\in\mathcal{R}$
The above model is a single-step variant of [18]. The original model is a multi-step planning problem, but we use its single step variant to ensure uniform modeling assumptions. Unlike [3], this model also has the travel time feasibility embedded in the last constraints:
$a_{ij}=\begin{cases}0~&{\mathcal{T}_{ij}\leq\mathcal{H}}\\ 1~&{\mathcal{T}_{ij}>\mathcal{H}}\end{cases}$ (33)
Similarly, for the allocation decision variable. A taxi in region $j\in\mathcal{R}$ can only be allocated to a passenger in region $i\in\mathcal{R}$ if the travel time is below a certain threshold $w$.
$b_{ij}=\begin{cases}0~&{\mathcal{T}_{ij}\leq w}\\ 1~&{\mathcal{T}_{ij}>w}\end{cases}$ (34)
For this study which uses a node-based approach, we only allow the drivers to allocate a driver to a passenger if the driver and passenger are in the same region (at same node). So $w$ is set to zero because we assume $\mathcal{T}_{ii}=0$, so only $\mathbf{Y_{ii}}$ is allowed to be nonzero.
Problem IV.6.
Model used for conversion of aggregate-level decisions of IV.4 and IV.5 to driver-level decisions.
$\max_{\mathbf{x}}$ $C$ (35)
s.t. $\sum_{j}\mathbf{x_{cj}}\leq 1\hskip 14.22636pt\forall c\in\mathcal{C},j\in\mathcal{R}$
$\sum_{c\in\mathcal{C}_{i}}\mathbf{x_{cj}}=\mathbf{X^{*}_{ij}}\hskip 14.22636pt\forall i\in\mathcal{R}$
Given the optimal rebalancing flows $\mathbf{X^{*}_{ij}}$ from IV.4 or IV.5, driver-level decisions are calculated using the above model, where the objective is any constant value.
Next, we discuss the case studies that we have performed to analyze the performance of the proposed AAVR model. In the first three case studies, we study different variants of two-station problem to elaborate the difference in recommendations by different models illustrating why adherence modeling is critical to ensure an effective vehicle rebalancing model in the context of taxi systems. Next, we illustrate the network level performance of the proposed model and benchmarks.
IV-D1 Case Study 1: Two-Station Problem - Impact of Driver Confidence
In the first case study, we first discuss the underlying logic of the rebalancing decisions generated by the vehicle rebalancing models in a setting where the drivers have a choice to accept or reject the system recommendation. Specifically, we use a two-station setting, say station A and B. Let there be 1000 taxis waiting at station A and 100 pickup requests at station B. The rebalancing system is fully aware of the supply and demand distribution. Further, each driver has a 50% confidence in system, i.e, $\mu(c)=0.5$. If a taxi driver is not recommended to go to station B or recommended to stay at station A, they stay at station A, i.e., $L_{cA}=1.0,L_{cB}=0.0$. If a driver accepts to move to station B, he is assumed to be able to reach it within the planning horizon. The rebalancing model has to decide how many and which taxi drivers should be recommended to go to station B. Logically, the recommender system must recommend 200 drivers to go to station B, so that in expectation 100 drivers move to station B; that would lead to all the demand being served without any waiting drivers. However, the baseline models are not designed considering the compliance uncertainties. Indeed, it is obvious to also note that since none of the baseline models consider the driver confidence, they must recommend the same number of drivers irrespective of the confidence levels of each driver assuming everything else remains the same.
The figure in the upper left of Fig. 9 shows the number of taxi drivers recommended to move to station B. Model B4 recommends 100 drivers to go to station B which in a setting where the confidence-level of all the drivers are completely adherent would be the optimal response if the objective is to serve all the passengers while ensuring that no driver remains idle. However, since the confidence level of each driver is 50%, only 50 out of 100 drivers accept the recommendation as shown in the top right figure in Fig. 9. This leads to 50 passenger being served with no driver remaining idle at station B. In contrast, B2 and B3 recommend all the 1000 drivers to move to station B. This is because both of these models try to match the supply proportion to the demand one. Since the demand proportion between stations A and B is 0: 1, that is, all passenger requests are at station B, all 1000 drivers are recommended to move to station B. In both of the cases B2, B3, all the demand is served by excessive number of drivers remain idle at the station B. The authors in B2 [6] do not discuss how to set the optimal value of the parameter $\beta$ to control the supply in such uncertain compliance setting. B1 uses a parameter $\rho$ to control the over-supply of the region. Additionally, they use a heuristic to not dispatch the exact number of taxis as the demand but rather a higher value depending on the travel time between the two stations. This leads to 500 recommendations when the travel time between station A and B is set to 4 mins and planning horizon is 5 mins. This again leads to fully satisfying the demand but 150 drivers remain idle. Without this heuristic, the model recommends 100 drivers $\rho=1$. So, the model has a flexibility to increase or decrease the flow. However, the authors in [7] also do not discuss how optimal $\rho$ can be selected in such uncertain compliance setting.
Finally, the AAVR model recommends 200 drivers to move to station B, out of which on average 100 drivers accept the recommendation, thus leading to all the demand being fulfilled and no driver remains idle at station B.
Remark 1.
The vehicle rebalancing models that are designed to dispatch the taxis by matching them with the demand such as B4, lead to sub-optimal demand fulfillment especially when confidence among drivers is low. In contrast, models such as B2 and B3 that do not control over-supply of the drivers lead to idle standing drivers, thus causing distrust among the drivers regarding the recommender system. A confidence-aware model such as the AAVR model is more robust to such compliance uncertainties.
Figure 9: Case Study 1: Two-Station Problem - Impact of Driver Confidence on Served Demand and Unallocated Drivers.
IV-D2 Case Study 2: Two-Station Problem - Impact of Driver Preference
In this case study, we investigate the impact of being aware of a taxi driver’s preference to reposition, i.e., the probability of moving to station B if he doesn’t receive any recommendation from the system, or rejects the system recommendation to stay in Station A. In this study, the drivers have a 50% confidence in the recommender system. Unlike case study 1, if the driver did not receive a recommendation or rejects the system’s recommendation to stay at station A, there is a 0.5 probability of the driver moving to station B, i.e., $L_{cA}=0.5,L_{cB}=0.5$.
All baseline models generate the same recommendations because there is no awareness of the driver’s confidence and preference. The situation among the drivers is further aggravated because in addition to the recommended drivers, some drivers reposition to station B on their own, thus leading to more competition for allocation on arriving at station B. The proposed model that is aware of the driver preference interestingly generates no recommendation, anticipating sufficient drivers repositioning on their own. Note that in this case, even if some drivers remain unallocated yet this does not impact the driver’s confidence on the recommender system because the repositioning decision was driver’s own.
Remark 2.
Information about driver preference is critical in designing an effective adherence-based rebalancing model. In cases such as above, the information about the driver preference may lead to fewer recommendations while achieving the objectives of higher served demands without risking the ineffective recommendations.
Figure 10: Case Study 2: Two-Station Problem - Impact of Driver Preference on Served Demand and Unallocated Drivers.
IV-D3 Case Study 3: Two-Station Problem - Driver Selection
In this case study, instead of each driver having a 0.5 confidence in the system’s recommendation, we initialize the driver confidence uniformly between 0 to 1. The driver preference is the same as in case study 1. As seen in Fig. 11 the baseline models as expected generate the same rebalancing decision, However compared to case study 1, in this case the proposed model recommends less than 200 drivers. This is further elaborated by the Fig. 12 which shows that the proposed model unlike the baseline models that are driver confidence agnostic, chooses the drivers with higher confidence in the recommender system while maintaining close to 100 driver in expectation. In this case, the proposed model, and baselines B1, B2, B3 are able to serve all the demands. However, no driver remains unallocated at station B for the proposed model, unlike that for baseline models B1, B2, and B3.
Figure 11: Case Study 3: Two-Station Problem - Impact of random confidence level on driver selection.
Figure 12: Case Study 3: Two-Station Problem - Distribution of the confidence level of the selected drivers.
Remark 3.
Although selecting drivers with higher confidence in the recommender system is desirable from the perspective of the platform, since it increases the likelihood of compliance with the rebalancing recommendations, this preference may introduce a bias that is not always ideal. It is important to note that this represents only a single moment in time. In subsequent case studies, we demonstrate that the proposed model not only accelerates the growth of driver confidence through repeated recommendations, but also improves both platform-centric and passenger-centric performance metrics.
TABLE III: Percentage Improvement Over Baselines B1–B4 Across Different Scenarios
Metric Optimistic Neutral Pessimistic
B1 B2 B3 B4 Avg B1 B2 B3 B4 Avg B1 B2 B3 B4 Avg
Served Demand 9.68 29.26 42.10 28.64 27.42 9.21 31.42 38.69 38.10 29.36 5.00 20.79 37.09 50.78 28.42
Waiting Time -11.49 -24.98 -26.25 -36.83 -24.89 -7.16 -21.79 -22.34 -31.41 -20.68 -13.14 -19.44 -22.98 -34.81 -22.59
Platform Earnings 9.42 27.78 42.62 29.97 27.45 8.99 30.26 38.79 39.75 29.45 5.07 19.77 37.51 53.55 28.97
Driver Profit 6.89 24.93 41.09 25.37 24.57 7.24 28.50 37.80 36.33 27.47 3.05 17.64 36.78 49.55 26.76
Figure 13: Comparison of proposed model with other baselines under optimistic, neutral, and pessimistic scenarios across four performance metrics.
IV-D4 Case Study 4: Network Level Problem
In this case study we do a complete network-level testing of the proposed model in comparison with the baseline models using Manhattan taxi network. The vehicle rebalancing model is executed after every 5 mins, considering the state of the drivers as well forecasted demand in the planning horizon. It is assumed that the demand arrives at the beginning of the time period and stays till the end of time period and drops out at the end of the time period if it remains un-served. First batch of the allocation happens at the beginning of the time period after which the remaining taxis are rebalanced to the nearby stations that can be reached within the planning horizon. Upon arrival the demands at the destination can be served by the taxis on the first come first serve basis. When a taxi is recommended to some station, the driver may accept or reject the recommendation based on the confidence level of the driver. We use 3000 taxis in simulation similar to baseline B1 [7] which leads to almost complete demand fulfillment in absolute adherence setting. All the taxi drivers start with a 50% confidence in the recommender system, suggesting unbiased opinion. The confidence evolves after each recommendation using the Beta-Bernoulli Thompson Sampling based dynamical system as explained in the Driver Confidence section II-B. To evaluate the performance of the proposed model as compared to the baseline models we use four metrics: i) Average Served Demand: This is defined as the total number of passenger requests served on average for a time period of 5 mins. ii) Average Waiting Time: This is defined as the time the passengers had to wait on average for allocation. The max wait time is 5 minutes beyond which the passenger leaves the system, iii) Average Platform Earning: This is defined as the total earnings by the platform. We assume 20% of each ride-share goes to the platform. iv) Average Driver Profit: This is defined as the difference between the ride fare after deduction by the platform and travel cost.
Table III reports the percentage improvement of the Proposed Model over baselines B1–B4 under three demand scenarios: Optimistic, Neutral, and Pessimistic. For Served Demand, the average improvements are 27.42%, 29.36%, and 28.42% respectively across the three scenarios. In terms of Waiting Time, the model achieves average reductions of 24.89%, 20.68%, and 22.59%. For Platform Earnings, the improvements are 27.45%, 29.45%, and 28.97%, while Driver Profit increases by 24.57%, 27.47%, and 26.76%. The gains remain consistently high across all metrics and scenarios, with the largest margins often observed in the pessimistic case. Among the baselines, B1 [7] emerges as the next-best performer across most metrics and scenarios. The empirical evidence shows that the heuristic used by [7] to bias the recommendations towards peak demand areas and recommending more vehicle than the predicted demand works well in this setting with adherence uncertainty, even though the model is not explicitly designed for it.
The results across all scenarios and fleet sizes indicate that the proposed model consistently improves Served Demand, Platform Earnings, and Driver Profit, while also reducing Waiting Time. In the Neutral scenario, Served Demand increased from 28.09% to 29.83% as fleet size grew from 2000 to 4000, while Waiting Time showed a significant reduction from 10.21-10.21- 10.21% to 47.07-47.07- 47.07%. Platform Earnings and Driver Profit also showed stable improvements, peaking at 29.94% and 27.47%, respectively. In the Optimistic scenario, the trend was similar, with Served Demand improving from 26.48% to 29.40%, and Waiting Time reducing sharply from 10.50-10.50- 10.50% to 54.63-54.63- 54.63%. Platform Earnings and Driver Profit reached up to 29.37% and 25.31%, respectively. The Pessimistic scenario yielded the most notable gains, particularly in Served Demand and Platform Earnings, which reached 37.73% and 38.33% at the 4000 fleet size. Despite the challenging conditions, Waiting Time was still significantly reduced from 8.05-8.05- 8.05% to 37.14-37.14- 37.14%. Overall, the results highlight the robustness of the model in improving key performance metrics across different demand scenarios and fleet sizes.